OpenStudy (anonymous):

Two positive numbers have a difference of 17. The larger number is one more than twice the smaller. Find the two numbers.

OpenStudy (radar):

Use algebra: Let x = larger positive number Let y = the other positive number x - y = 17 Given by problem. x = 2y + 1 also given by the problem. Substitute this value of x in first equation 2y + 1 - y = 17 Can you take it from there.

OpenStudy (anonymous):

yes thank you so much!

OpenStudy (radar):

No problem, good luck with written problems, they do sprain the brain.

OpenStudy (anonymous):

lol i think my brain has already broke!

OpenStudy (radar):

It helps to verify (time permitting) by testing the results. In this case I came up with y =16 and x = 33. First test does x-y=17 33-16= 17 checks. Does x = 2(16) + 1 x=32 + 1 = 33 Checks, has to be a correct solution.